# The Conspiracy Money Machine: Uncovering

### v2

@@ 7 Mitigation
The Channel Checker Bot. It is a Telegram Bot implemented in Python3. It receives a channel name from the user and provides detailed information about that channel. The bot uses a database containing channels identified during our research and information such as questionable content, advertised products, and dubious fundraising campaigns. When a user inputs a channel name, the bot searches the database for a match. If the channel is found, the bot provides the user with the existing information from the database. Instead, if the channel is not in the database, the bot initiates a headless Telegram client, joins the channel via Telegram APIs, and retrieves all shared content. Then, it extracts all shared links and checks them against the conspiracy resource dataset and the monetization links collected during this research. After the analysis, the bot updates the database with the new channel and its associated information. Finally, the bot presents the analysis result to the user. Fig. 6 displays an instance of a chat with the Channel Checker Bot. Therefore, the Channel Checker Bot serves a dual purpose: It provides users with detailed information about channels and continuously updates the database with new conspiracy channels.
@@
ConspiracyAlert: A Browser Plug-in. Although the Channel Checker Bot serves as a warning system for Telegram users navigating the platform, it does not help users who browse the web and land on questionable products or crowdfunding campaigns. To address this gap, we developed the ConspiracyAlert browser plug-in. This plug-in, compatible with Chrome and Firefox and written in TypeScript, monitors the user’s web navigation. Once a webpage is fully loaded, the plug-in captures the URL and transmits it to a remote server. Then, the server checks the URL against both the Conspiracy Resource dataset and the URLs dataset identified in Sec. 4.2.2. If the server finds a match, it generates an informative message, which the plug-in displays to the user.
The source code for both tools and additional technical implementation details are publicly available on GitHub [14]. Notably, we designed both tools to maintain user privacy by not retaining user information (e.g., IP addresses) or details that could later associate a request with a user (e.g., timestamps, Telegram client versions, or browsing history). The system only stores information about the analyzed channels.
@@ 8 Limitations
As mentioned in Sec. 4.2, we build the Conspiracy Resources Dataset gathering information from previous work focused on conspiracy theories. However, most of the scientific literature focuses on analyzing English content. This limitation could introduce a bias, as conspiracy communities operating in other languages, such as Russian or Indian, might evade detection due to their use of non-English sources. Moreover, other platforms not considered in our study (e.g., Parler [17]) are known to host conspiracy-related content. Unfortunately, we could not find works providing conspiracy-labeled datasets suitable for our study.
This study primarily utilizes link analysis to uncover monetization techniques and estimate the amount of money raised. However, this approach does not provide insight into monetization campaigns conducted through media content. For example, we observed various images shared on the channels promoting self-published books, paid events (e.g., conferences), or training courses conducted by self-proclaimed gurus. In Sec. 5.2, we analyzed campaigns that collected more than \$200,000 and found several that, despite being endorsed by conspiracy channels, do not show evidence of being connected to conspiracy theories. This suggests that other similar campaigns may exist among those we considered in this study. Understanding why these campaigns are endorsed by conspiracy channels raises intriguing questions: Are they shared only to support the conspiracy narratives? Are their topics intentionally unrelated to conspiracy theories to appeal to a broader audience? Or are the administrators paid to promote these campaigns?
@@ 10 Conclusion and future work
In this work, we focused on understanding and quantifying how conspiracy theories raise funds by exploiting Telegram. We started by identifying the conspiracy theory-related channels, analyzing a novel dataset we built by collecting previously validated resources from an extensive literature review that we publicly release [15].
This study revealed the alarming finding that almost 15% of all Telegram channels in the TGDataset (17,829 channels) are linked to conspiracy theories. Then, we discover that conspiracy theory-related channels actively seek to profit from their subscribers. We categorize all the diverse monetization strategies we find in our dataset and dive into the analysis of the three most common. Our study shows that conspiracy theories raised funds for about \$66 million by arranging crowdfunding campaigns.
@@ Appendix A ChatGPT prompt
The input provided is a text containing one or more URLs pointing to donation or crowdfunding platform. Classify the message with 1 if the message is asking directly or indirectly for donations, or to support a project. Otherwise, return 0 if the message is discrediting or asking to report the recipient of the donation or the project.

### v3

@@ 7 Mitigation
The Channel Checker Bot. It is a Telegram Bot implemented in Python. It receives a channel name from the user and provides detailed information about that channel. The bot uses a database containing channels identified during our research and information such as questionable content, advertised products, and dubious fundraising campaigns. When a user inputs a channel name, the bot searches the database for a match. If the channel is found, the bot provides the user with the existing information from the database. Instead, if the channel is not in the database, the bot initiates a headless Telegram client, joins the channel via Telegram APIs, and retrieves all shared content. Then, it extracts all shared links and checks them against the conspiracy resource dataset and the monetization links collected during this research. After the analysis, the bot updates the database with the new channel and its associated information. Finally, the bot presents the analysis result to the user. Fig. 6 displays an instance of a chat with the Channel Checker Bot. Therefore, the Channel Checker Bot serves a dual purpose: It provides users with detailed information about channels and continuously updates the database with new conspiracy channels.
@@
ConspiracyAlert: A Browser Plug-in. Although the Channel Checker Bot serves as a warning system for Telegram users navigating the platform, it does not help users who browse the web and land on questionable products or crowdfunding campaigns. To address this gap, we developed the ConspiracyAlert browser plug-in. This plug-in, compatible with Chrome and written in TypeScript, monitors the user’s web navigation. Once a webpage is fully loaded, the plug-in captures the URL and transmits it to a remote server. Then, the server checks the URL against both the Conspiracy Resource dataset and the URLs dataset identified in Sec. 4.2.2. If the server finds a match, it generates an informative message, which the plug-in displays to the user.
The source code for both tools and additional technical implementation details are publicly available on GitHub [79, 81]. Notably, we designed both tools to maintain user privacy by not retaining user information (e.g., IP addresses) or details that could later associate a request with a user (e.g., timestamps, Telegram client versions, or browsing history). The system only stores information about the analyzed channels.
@@ 8 Limitations
As mentioned in Sec. 4.2, we build the Conspiracy Resource Dataset gathering information from previous work focused on conspiracy theories. However, most of the scientific literature focuses on analyzing English content. This limitation could introduce bias, as conspiracy communities operating in other languages, such as Russian or Indian, might evade detection due to their use of non-English sources. Moreover, other platforms not considered in our study (e.g., Parler [13]) are known to host conspiracy-related content. Unfortunately, we could not find works providing conspiracy-labeled datasets suitable for our study.
This study primarily utilizes link analysis to uncover monetization techniques and estimate the amount of money raised. This approach does not provide insight into monetization campaigns conducted through media content. For example, we observed various images shared on the channels promoting self-published books, paid events (e.g., conferences), or training courses conducted by self-proclaimed gurus. In Sec. 5.2, we analyzed campaigns that collected more than \$200,000 and found several that, despite being endorsed by conspiracy channels, do not show evidence of being connected to conspiracy theories. This suggests that other similar campaigns may exist among those we considered in this study. Understanding why these campaigns are endorsed by conspiracy channels raises intriguing questions: Are they shared only to support the conspiracy narratives? Are their topics intentionally unrelated to conspiracy theories to appeal to a broader audience? Are administrators paid to promote these campaigns?
@@ 10 Conclusion and future work
In this work, we focused on understanding and quantifying how conspiracy theories raise funds by exploiting Telegram. We started by identifying the conspiracy theory-related channels, analyzing a novel dataset we built by collecting previously validated resources from an extensive literature review that we publicly release [80]. This study revealed the alarming finding that almost 15% of all Telegram channels in the TGDataset (17,829 channels) are linked to conspiracy theories. Then, we discover that conspiracy theory-related channels actively seek to profit from their subscribers. We categorize all the diverse monetization strategies we find in our dataset and dive into the analysis of the three most common. Our study shows that conspiracy theories raised funds for about \$71 million by arranging crowdfunding campaigns.
@@
Acknowledgments
This work has been partially funded by projects: MUR National Recovery and Resilience Plan, SERICS (PE00000014); and ST3P (B83C24003210001) under the "Young Researchers 2024-SoE" Program funded by the Italian Ministry of University and Research (MUR). We express our gratitude to Lucia Fores for her contributions to the Channel Checker Bot and the Conspiracy Alert browser plugin.
@@ Appendix A ChatGPT prompt
The input provided is a text containing one or more URLs pointing to donation or crowdfunding platform. Classify the message with 1 if it is asking directly or indirectly for donations, or to support a project. Otherwise, return 0 if the message is discrediting or asking to report the recipient of the donation or the project.